OpenStudy (anonymous):

Find the equation of the curve that passes through the point (1,3) and has a slope of y/x^2 at the point (1,3). i used y-y1=m(x-x1) and ended up with y=3x, however that is linear but the problem said it was a curve. Does that seem right?

myininaya (myininaya):

So x=1,y=3 what is the slope?

myininaya (myininaya):

it says the general form for the slope is y/x^2

myininaya (myininaya):

So plug in and evaluate

myininaya (myininaya):

have you replaced y with 3 and x with 1 yet?

myininaya (myininaya):

Yes i think they are looking for a tangent line to the curve at (1,3) tangent lines are linear and have the form you speak of

myininaya (myininaya):

wait unless ...

myininaya (myininaya):

maybe it does actually mean find the equation to the curve is this calculus 2?

myininaya (myininaya):

sweetlove333 are you there?

myininaya (myininaya):

So we are given y'=y/x^2 We can use separation of varaibles to find the equation of the curve y

myininaya (myininaya):

then we have the point (1,3) so we can find the constant C after integrating

OpenStudy (anonymous):

i'm in ap calc ab, can you show the steps?

myininaya (myininaya):

have you talk about integration?

OpenStudy (anonymous):

yeah, do you think i should integrate the y/x^2 equation? and then plug in the values for x,y to find c?

myininaya (myininaya):

First y'=y/x^2 We need to separate the variables put all of our y's on one side and all of our x's on the other

myininaya (myininaya):

\[\frac{dy}{dx}=\frac{y}{x^2}\]

myininaya (myininaya):

So we need to multiply 1/y on both sides given us \[\frac{1}{y} \frac{dy}{dx}=\frac{1}{x^2}\] And then we need to multiply dx on both sides given us \[\frac{1}{y} dy =\frac{1}{x^2} dx\] Now this gives you permission to integrate both sides Can you do that part? Tell me what you get...

OpenStudy (anonymous):

after integrating, I got ln absolute value y = (but how do you antiderive( dx/x^2)?

OpenStudy (anonymous):

can i use the trig fcn of arctan?

myininaya (myininaya):

\[\text{ recall from algebra you can rewrite } \frac{1}{x^2} \text{ as } x^{-2}\]

myininaya (myininaya):

Do you know the power rule for integration?

OpenStudy (anonymous):

so it would -x^-1?

myininaya (myininaya):

:)

myininaya (myininaya):

ok and then +c on of of the sides

myininaya (myininaya):

one of*

myininaya (myininaya):

\[\ln|y|=\frac{-1}{x}+C\]

myininaya (myininaya):

right?

OpenStudy (anonymous):

okay that's what i got, then do i plug in the points and figure out c?

myininaya (myininaya):

So we can find C since we are given a point on the curve (1,3)

myininaya (myininaya):

lol yes sweetlove great! :)

OpenStudy (anonymous):

thanks!, and after finding c, what equation should i plug c into then?

OpenStudy (anonymous):

would it be y=lny+ x^-1 +c? and substitute the c i found?

myininaya (myininaya):

You are looking for the curve (we just found the general form for the curve) so when you find C replace C in the equation and you will have your curve y

OpenStudy (anonymous):

would that be the equation of the curve?

myininaya (myininaya):

where that extra y come from

myininaya (myininaya):

\[\ln|y|=\frac{-1}{x}+C \] This is the general form for the curve

OpenStudy (anonymous):

okay, thanks a lot! =)

myininaya (myininaya):

NP What did you get for C?

OpenStudy (anonymous):

for c i got 2.0986

myininaya (myininaya):

Are you allowed to approximate?

OpenStudy (anonymous):

i hope so=P

myininaya (myininaya):

\[\ln|3|=\frac{-1}{1}+C\] So this is what you have where you replace x with 1 and y with 3

myininaya (myininaya):

now this means we have ln(3)=-1+C

myininaya (myininaya):

So adding 1 on both sides will isolate C

myininaya (myininaya):

ln(3)+1=C C=ln(3)+1=ln(3)+ln(e)=ln(3e) <---you don't have to write this way (I just think it looks prerrier; you can leave it as ln(3)+1)

myininaya (myininaya):

So you can write the equation of the curve like this \[\ln|y|=\frac{-1}{x}+(\ln(3)+1)\] OR (my favorite) \[\ln|y|=\frac{-1}{x}+\ln(3e)\]

OpenStudy (anonymous):

okay, that was very helpful! ^______^ thanks!

myininaya (myininaya):

:) great sweetlove I actually like this problem

myininaya (myininaya):

You know after I read your problem again it does not say anything about the tangent line so I'm sure that that is our answer above